OpenStudy (anonymous):

Ax^2+Cy^2+F=0 A,C,F don't equal 0 A & C are of the same sign and F is of the opposite sigh what is the equation of the circle when the center (0,0) if A equals C

OpenStudy (anonymous):

if \(A=C\) then it is \[Ax^2+Ax=-F\]

OpenStudy (anonymous):

oops i meant \[Ax^2+Ay^2=-F\]

OpenStudy (anonymous):

or if you prefer \[x^2+y^2=-\frac{F}{A}\]